# VLCE: A Knowledge-Enhanced Framework for Image Description in Disaster Assessment

### v3

@@ Abstract
Immediate damage assessment is essential after natural catastrophes; yet, conventional hand evaluation techniques are sluggish and perilous. Although satellite and unmanned aerial vehicle (UAV) photos offer extensive perspectives of impacted regions, current computer vision methodologies generally yield just classification labels or segmentation masks, so constraining their capacity to deliver a thorough situational comprehension. We introduce the Vision Language Caption Enhancer (VLCE), a multimodal system designed to produce comprehensive, contextually-informed explanations of disaster imagery. VLCE employs a dual-architecture approach: a CNN-LSTM model with a ResNet50 backbone pretrained on EuroSat satellite imagery for the xBD dataset, and a Vision Transformer (ViT) model pretrained on UAV pictures for the RescueNet dataset. Both systems utilize external semantic knowledge from ConceptNet and WordNet to expand vocabulary coverage and improve description accuracy. We assess VLCE in comparison to leading vision-language models (LLaVA and QwenVL) utilizing CLIPScore for semantic alignment and InfoMetIC for caption informativeness. Experimental findings indicate that VLCE markedly surpasses baseline models, attaining a maximum of 95.33% on InfoMetIC while preserving competitive semantic alignment. Our dual-architecture system demonstrates significant potential for improving disaster damage assessment by automating the production of actionable, information-dense descriptions from satellite and drone photos.
@@ 1 Introduction
Global populations are particularly vulnerable to natural catastrophes, since floods, earthquakes, hurricanes, and wildfires cause extensive harm to ecosystems, infrastructure, and human lives all across the world. In these circumstances, quick damage assessment is essential for directing long-term restoration activities, assisting with search and rescue missions, and assigning emergency resources. However, in the immediate aftermath of disasters, typical manual assessment procedures are sometimes impracticable, slow, and dangerous.
@@
Example (Disaster Scenario):
Figure 1: RescueNet disaster image of post-hurricane devastation contrasting basic visual description with detailed automatic caption generation.
Without caption: “A satellite image depicts a community from an aerial perspective, revealing multiple dwellings and streets”
With caption: “The image shows the aftermath of Hurricane Michael, which inflicted major damage to infrastructure and the environment. Roads and streets are cluttered with debris from destroyed structures, such as fallen trees and scattered items. Trees in the region have suffered varied degrees of damage, with some seeming damaged or totally fallen. Buildings show different degrees of destruction, ranging from comparatively intact ones to those with obvious traces of catastrophic damage. The entire scene depicts the ongoing recovery and rebuilding activities in the aftermath of the catastrophe.”
@@
To overcome these limitations, we propose VLCE (Vision Language Caption Enhancer), a specialized image captioning framework that enhances caption generation by integrating knowledge for the analysis of disaster scenarios. This approach combines visual features with external semantic knowledge from ConceptNet and WordNet to improve semantic understanding and produce descriptive, contextually relevant captions. Our model utilizes cross-modal attention mechanisms within tailored architectures to connect raw visual signals with significant, domain-specific textual descriptions. This framework specifically targets the generation of descriptions for satellite and drone imagery in disaster-affected regions, thereby addressing a notable deficiency in automated disaster assessment capabilities.

### v4

@@ Abstract
The processes of classification and segmentation utilizing artificial intelligence play a vital role in the automation of disaster assessments. However, contemporary VLMs produce details that are inadequately aligned with the objectives of disaster assessment, primarily due to their deficiency in domain knowledge and the absence of a more refined descriptive process. This research presents the Vision Language Caption Enhancer (VLCE), a dedicated multimodal framework aimed at integrating external semantic knowledge from ConceptNet and WordNet to improve the captioning process. The objective is to produce disaster-specific descriptions that effectively convert raw visual data into actionable intelligence. VLCE utilizes two separate architectures: a CNN-LSTM model that incorporates a ResNet50 backbone, pretrained on EuroSat for satellite imagery (xBD dataset), and a Vision Transformer developed for UAV imagery (RescueNet dataset). In various architectural frameworks and datasets, VLCE exhibits a consistent advantage over baseline models such as LLaVA and QwenVL. Our optimal configuration reaches an impressive 95.33% on InfoMetIC for UAV imagery while also demonstrating strong performance across satellite imagery. The proposed framework signifies a significant transition from basic visual classification to the generation of comprehensive situational intelligence, demonstrating immediate applicability for implementation in real-time disaster assessment systems.
@@ 1 Introduction
Global populations are particularly vulnerable to natural catastrophes, since floods, earthquakes, hurricanes, and wildfires cause extensive harm to ecosystems, infrastructure, and human lives all across the world. In these circumstances, quick damage assessment is essential for directing long-term restoration activities, assisting with search and rescue missions, and assigning emergency resources. However, typical manual assessment procedures in the aftermath of the disasters are often impracticable, time-consuming, and constrained by safety concerns.
@@
The Value of Contextual Captioning [Figure 1]:
Figure 1: Post-hurricane disaster imagery from the RescueNet dataset.
Without knowledge graph caption: “A satellite image depicts a community from an aerial perspective, revealing multiple dwellings and streets”
With knowledge graph caption: “The image shows the aftermath of Hurricane Michael, which inflicted major damage to infrastructure and the environment. Roads and streets are cluttered with debris from destroyed structures, such as fallen trees and scattered items. Trees in the region have suffered varied degrees of damage, with some seeming damaged or totally fallen. Buildings show different degrees of destruction, ranging from comparatively intact ones to those with obvious traces of catastrophic damage. The entire scene depicts the ongoing recovery and rebuilding activities in the aftermath of the catastrophe.”
@@
To address these limitations, we propose VLCE (Vision Language Caption Enhancer), a framework that enhances generated captions by integrating knowledge for the analysis of disaster scenarios. This approach combines visual features with external semantic knowledge from ConceptNet and WordNet to improve semantic understanding and produce descriptive, contextually relevant captions. Our model utilizes cross-modal attention mechanisms within tailored architectures to connect raw visual signals with significant, domain-specific textual descriptions. This framework specifically targets the generation of descriptions for satellite and drone imagery in disaster-affected regions, thereby addressing a notable deficiency in automated disaster assessment capabilities.